OpenStudy (anonymous):

What is the value of z in the equation 4(2z + 3) = 12?

iYuko (iyuko):

First distribute the 4 to make (8z+12)=12

OpenStudy (anonymous):

Alright.

iYuko (iyuko):

Then subtract 12 from both sides

iYuko (iyuko):

Wait a minute

OpenStudy (anonymous):

Sooo 8z-12?

iYuko (iyuko):

Divide by 8 sorry

OpenStudy (anonymous):

Okay soo 8z divided by 8 and 12 divided by 8?

iYuko (iyuko):

to get z+12=1.5

OpenStudy (anonymous):

So z=1.5

iYuko (iyuko):

then subtract 12 from both sides

iYuko (iyuko):

z=???

OpenStudy (anonymous):

12-12=0

iYuko (iyuko):

1.5-12

OpenStudy (anonymous):

-10.5

OpenStudy (anonymous):

Soo z=-10.5?